75% of first order reaction is completed in 30 minutes.what is the time required for 93.75% of the reaction in minutes

Grade:12th pass

1 Answers

Dharma Teja
7 years ago
as this is a first order reaction
75=50+25=30min
50+25+12.5=87.5=60min
50+25+12.5+6.25=93.75=120min
click approve if you like the answer